OpenStudy (anonymous):

Simplify & show work: x-1/3 = x+1/4

OpenStudy (anonymous):

This gives a false answer. By adding 1/3 to both sides then subtracting x from both sides you are told that ... 0 = 7/12

OpenStudy (anonymous):

We both know this is not true!

OpenStudy (anonymous):

There is an answer to it.You need to find the common denominator or something.

OpenStudy (anonymous):

I know the aswer is 7. I don't know how to get the answer though.

OpenStudy (anonymous):

When you find it OMGI let me know :)

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

Irene just think logically. You are told if you add something to x it is the same result as taking something away from x. See it like this I am thinking of a number that when you add it to a euro it is the same as taking it away from a euro. The onlt thing it can be is 0

OpenStudy (anonymous):

...and even this does not satisfy your equation

OpenStudy (anonymous):

I still believe there is an answer... lol. Maybe the common denominator is 12 since the denominators need to be multiplied.

OpenStudy (anonymous):

Sorry Irene my analogy was rubbish! If you add 1/4 to a euro dollar whatever, it is the same as subtracting 1/3 from the same euro. It does not add up, so to speak!

OpenStudy (anonymous):

I am confused.

OpenStudy (anonymous):

I will do it line by line

OpenStudy (anonymous):

there is no answer to it the given equation is wrong........................` x gets cancelled

OpenStudy (anonymous):

believe me............................... i am sure i am studying in class 8(dont worry)

OpenStudy (anonymous):

x-1/3 = x+1/4 add 1/3 to both sides x = x + 1/4 + 1/3 now subtracrt x from both sides so separating variables and constants 0 = 1/4 + 1/3 0= 7/12 Are you happy that 7/12 is equal to 0?

OpenStudy (anonymous):

Okay, thank you then!

OpenStudy (anonymous):

welcome.............................

OpenStudy (anonymous):

Wait, did you read it wrong? I meant x-1 OVER 3 = x + 1 OVER 4.

OpenStudy (anonymous):

OK it does not read that way

OpenStudy (anonymous):

In this case x = 7

OpenStudy (anonymous):

How does it equal 7?

OpenStudy (anonymous):

then yhe answer is (x-1)4=(x+1)3 4x-4=3x+3 x=7 happy..................

OpenStudy (anonymous):

So you cross multiply in the beginning? And thank you, i'm happy now!

OpenStudy (anonymous):

welcome............yes,u have to cross multiply the denominators................

OpenStudy (anonymous):

(x - 1)/3 = (x + 1)/4 multiply both sides by 3 x-1 = 3 (x + 1)/4 multiply both sides by 4 4 (x-1)= 3 (x + 1 Expand bracketrs 4x-4 = = 3x + 3 Subtract 3x from both sides x-4 = 3 add 4 to both sides x = 7

OpenStudy (anonymous):

Thank you both of you! I understand it now! :)

OpenStudy (anonymous):

welcome once again

OpenStudy (anonymous):

Great, well done for persevering Irene